OpenStudy (anonymous):

Combine the rational expressions and simplify: 3x/9x + 8/x

OpenStudy (lalaly):

x/3x + 24/3x = x+24/3x

OpenStudy (anonymous):

let us first simplify the first equation 3x/9x = x/3x so we have x/3x +8/x. Now to ad wee have to get a common denominator. So we multiply 8/x by 3/3 = 24/3x x/3x +24/3x = (x+24)/ 3x

OpenStudy (anonymous):

Ahh, ok, the y cancelled out.

OpenStudy (lalaly):

its not wrong

OpenStudy (anonymous):

I was wrong to call you wrong, I read it all wrong xC I'm sorry

OpenStudy (anonymous):

Did you mean for there to be a y?

OpenStudy (lalaly):

its ok :D:D

OpenStudy (anonymous):

Yes, there's a y in the equation. When I combine the problems I get x^2+24y/3xy. I know that's wrong. Do I need to simplify that to get my real answer?

OpenStudy (anonymous):

Well, you need to modify your problem statement. Perhaps you mistyped it.

OpenStudy (anonymous):

Ahh yes :( I did. I'm very sorry.... It's 3x/9y + 8/x